OpenStudy (anonymous):

What is the equation of the line that passes through the point (-2, 2) and has a slope of 5?? Please help, I'm so confused on how to do this :(

OpenStudy (rina.r):

y=mx+c so, y=2x+5 i think

OpenStudy (rina.r):

i am sure about 2x the gradient b coz in the bracket the first variable is always x which is -2 and then y which is 2 so yeah.

hartnn (hartnn):

use the slope point form \(y-y_1 = m (x-x_1 )\)

OpenStudy (rina.r):

i think the formula is y2-y1/x2-x1

hartnn (hartnn):

m = 5 x1 = -2 y1 = 2

OpenStudy (rina.r):